# Curriculum Direct Preference Optimization for Diffusion and Consistency Models

### v2

@@ A.3 More Qualitative Results
Figure 5: Qualitative results after fine-tuning with the LAION Aesthetic Predictor as the reward model. The fine-tuning alternatives are: DDPO, DPO and Curriculum DPO. Best viewed in color.
@@
In Figures 4 and 5, we present qualitative results after fine-tuning the models with HPSv2 and LAION Aesthetic Predictor as reward models. Fine-tuning for human preference generally results in generating images with more details for the LCM model. Curriculum DPO, in particular, produces better aesthetics for both foreground objects and the background. On the other hand, the SD results show a better alignment with the text prompt, Curriculum DPO being the only method that generates the ant on a bike displayed in the last column. Fine-tuning for improving the visual appeal returns in general, as expected, better aesthetics for the animals. However, Curriculum DPO returns several examples that look better, e.g. the camel in the sixth column and the dog in the third column.
@@
A.4 Human Evaluation Study
@@
Since there are four images for each prompt and an annotation form comprises 80 prompts, the number of images to be annotated in one form is 320. Each form was completed by four human evaluators, yielding a total of 1,280 annotations per experiment (form). Since we conducted the study on two generative models (SD and LCM) and two tasks (prompt alignment and aesthetics), the total number of collected annotations is 5,120.
The average time to complete the annotations for a single form is around 15-20 minutes. The four human annotators who agreed to complete the annotation forms are either close collaborators, family members or friends of the authors. They volunteered to perform the annotations for free. To make sure that the annotations are relevant, we computed the inter-rater agreement, obtaining a Kendall Tau correlation coefficient of $0.34$. This translates into $69.8\%$ of all image pairs being concordant among annotators.
A.5 Limitations
One limitation of our model is the introduction of additional hyperparameters, such as $B$ or $K$. These might require tuning in order to find the optimal values, which requires more computing power. However, in the ablation study from Section 4, we demonstrate that Curriculum DPO outperforms all baselines for multiple hyperparameter combinations. Therefore, suboptimal hyperparameter choices can still improve the generative models.
@@ A.5 Limitations
A.6 Broader Impacts
Generative models can be a valuable asset in many scenarios, ranging from boosting the productivity of creative tasks to being integrated in applications that are used on a daily basis (such as image restoration or super-resolution). Nevertheless, it might also represent a great source of fake data aiming for disinformation and impersonation, especially when the model is optimized to human preferences. In the recent years, an increase in deep fake materials flooded the Internet, spreading false information or even intending to steal sensitive information by posing as another entity or person.
While we strongly believe in the benefits of very capable generative models, we are aware of the potential risks. However, we can see that governments are working very closely with academia and industry on safely developing artificial intelligence, and thus observe and support the increasing focus on models that detect AI-generated content to mitigate the aforementioned risks.

### v3

@@ 7 Curriculum DPO for Diffusion Models
8 Importance of Consistency-DPO
Our work makes two contributions: Curriculum DPO and Consistency-DPO. While the novelty and importance of Curriculum DPO is more obvious, we consider that the significance of Consistency-DPO is not immediately observable. To this end, it is important to note that the Diffusion-DPO [72] approach cannot be directly applied to consistency models. The most direct modification is to substitute the noise estimation in the Diffusion-DPO objective with the consistency distillation loss used in consistency models. However, applying this modification directly breaks the consistency property required by these models and leads to poor results, as shown in Figure 4 and further discussed in Section 10.
We found two solutions for this problem. The first is to reintegrate the consistency distillation for both preferred and non-preferred samples as separate components within the optimization function, in addition to the Consistency-DPO loss (Eq. (6)). This method, however, introduces the need for additional hyperparameters to balance these terms, which represents a significant drawback because it requires extensive hyperparameter tuning.
The second solution, which we ultimately adopt in our study, is to ensure the initial estimation for the ODE’s starting point (the target in the consistency distillation loss) is a sample of the consistency model that undergoes fine-tuning. We accomplish this by replacing the Exponential Moving Average (EMA) model, that is typically used to get this estimation, with the pre-trained model from which we begin the fine-tuning process. This approach maintains the integrity of the consistency model’s properties throughout the training.
We thus conclude that adapting DPO to consistency models is not trivial, since the adaptation requires a deep understanding of the framework and strong knowledge about the use of gradients.
9 More Quantitative Results
Fine-Tuning Text Aesthetics Human
Strategy Alignment Preference
- 0.5246 5.6675 0.2673
DDPO 0.5317 5.6764 0.2717
DPO 0.5328 5.7593 0.2725
Curriculum DPO (ours) 0.5413 5.7998 0.2783
Table 4: Text alignment, aesthetic and human preference scores on Pick-a-Pic ($D_{3}$), obtained by the baseline (pre-trained) Stable Diffusion model versus the three fine-tuning strategies: DDPO, DPO and Curriculum DPO. The best scores are highlighted in bold.
Results on Pick-a-Pic. We report additional results for Stable Diffusion on 150,000 image pairs from Pick-a-Pic ($D_{3}$) in Table 4. In this scenario, the dataset already includes pairs of winning and losing images, so we only apply the reward models for the ranking described in Figure 1. The results reported on $D_{3}$ are consistent with those reported on ${D}_{1}$ and ${D}_{2}$, further highlighting the importance of curriculum learning.
Fine-Tuning Strategy LLaVA Phi-3
- 0.6804 0.6804
DDPO 0.7629 0.7602
DPO 0.7614 0.7643
Curriculum DPO (ours) 0.7703 0.7736
Table 5: Text alignment results on dataset $D_{1}$ by using two reward models (LLaVA and Phi-3) for DDPO, DPO and Curriculum DPO applied on Stable Diffusion. The best scores are highlighted in bold.
Results with different reward models. In Table 5, we compare text alignment results for two alternative reward models: LLaVA [33] and Phi-3 [1]. During training, we use a reward model to extract image descriptions and then measure their similarity to the original prompts to produce winning and losing image pairs. The same similarity scores also determine the ranking used by Curriculum DPO. These experiments further confirm the superiority of Curriculum DPO over DPO and DDPO, regardless of the employed reward model.
10 More Qualitative Results
@@
In Figures 5 and 6, we present qualitative results after fine-tuning the models with HPSv2 and LAION Aesthetics Predictor as reward models on $D_{1}$, respectively. Fine-tuning for human preference (Figure 5) generally results in generating images with more details for the LCM model. Curriculum DPO, in particular, produces better aesthetics for both foreground objects and the background. In contrast, the SD results show a better alignment with the text prompt, Curriculum DPO being the only method that generates the ant on a bike displayed in the last column. Fine-tuning for improving the visual appeal (Figure 6) returns in general, as expected, better aesthetics for the animals. However, Curriculum DPO returns several examples that look better, e.g. the camel in the sixth column and the dog in the third column.
Figure 6: Qualitative results after fine-tuning with the LAION Aesthetics Predictor as the reward model. The fine-tuning alternatives are: DDPO, DPO and Curriculum DPO. Best viewed in color.
(a) Varying the learning rate for Curriculum DPO on LCM.
(b) Varying the number of LCM generation steps for Curriculum DPO.
(c) The progression of the aesthetic reward function during training with the DDPO method.
Figure 7: Additional ablation results for Curriculum DPO applied on LCM are depicted in Figure 7(a) and 7(b). In Figure 7(c), we show the evolution of the reward score when training Stable Diffusion with DDPO. All experiments are carried out on DrawBench.
In Figure 4, we show qualitative results when fine-tuning the model for text alignment on $D_{2}$. In addition to the baseline, DPO, DDPO and Curriculum DPO results, we also include the images generated by a naive implementation of Consistency-DPO. This implementation refers to the most direct adaptation of Diffusion-DPO to consistency models. More precisely, we substitute the noise estimation in the Diffusion-DPO objective with the consistency distillation loss used in consistency models. However, applying this modification directly breaks the consistency property required by these models and leads to bad results, as illustrated in the 4th row of Figure 4.
@@
LCM - 0.5602 5.8038 0.2610
DDPO 0.5627 5.9488 0.2780
DPO 0.5639 5.9611 0.2783
Curriculum DPO (ours) 0.5654 6.0038 0.2793
Table 6: Text alignment, aesthetic and human preference scores obtained on the DrawBench dataset by the baseline (pre-trained) LCM versus the three fine-tuning strategies: DDPO, DPO and Curriculum DPO. The DDPO, DPO and Curriculum DPO methods use only 5 images per prompt during optimization. The best scores are highlighted in bold.
11 Additional Ablations
Aside from the ablation results presented in Figure 3 and Table 3 from the main article, there are a few other hyperparameters involved in the training process, such as the learning rate and the number of steps used in the multi-step generation of LCM. We performed additional ablation studies on the learning rate and the number of steps used by LCM, on the DrawBench dataset, using $M=5$ generated images per prompt. The results presented in Figures 7(a) and 7(b) demonstrate that, regardless of the chosen values, the outcomes consistently surpass the baseline (see Table 6). We emphasize that we did not try to tune these hyperparameters for Curriculum DPO to avoid overfitting in hyperparameter space. Moreover, we underline that some apparent hyperparameters directly depend on already ablated hyperparameters. For example, the hyperparameter $B$ (ablated in Figure 3(c)) is the only one that influences the minimum/maximum preference limits $\mathrm{L}_{k}$ and $\mathrm{R}_{k}$, which are computed in steps 3 and 4 of both Algorithm 1 and Algorithm 2. Note that the equations for $\mathrm{L}_{k}$ and $\mathrm{R}_{k}$ generate equally-sized batches, so the limits change only when we change the number of curriculum batches $B$. Therefore, ablating $\mathrm{L}_{k}$ and $\mathrm{R}_{k}$ is redundant.
In Figure 7(c), we present the evolution of the reward score during the Stable Diffusion training with DDPO. For DPO and Curriculum DPO, we did not preserve the reward curves, as these methods did not involve multiple queries to the reward models. Instead, they rely solely on the original example ranking throughout the entire training process. Thus, additional queries to the reward models are unnecessary for DPO and Curriculum DPO. This represents an advantage of these methods over DDPO.
12 Human Evaluation Study
@@
Since there are four images for each prompt and an annotation form comprises 80 prompts, the number of images to be annotated in one form is 320. Each form was completed by nine human evaluators, yielding a total of 2,880 annotations per experiment (form). Since we conducted the study on two generative models (SD and LCM) and two tasks (prompt alignment and aesthetics), the total number of collected annotations is 11,520.
The average time to complete the annotations for a single form is around 15-20 minutes. The nine human annotators who agreed to complete the annotation forms are either close collaborators, family members or friends of the authors. They volunteered to perform the annotations for free. To make sure that the annotations are relevant, we computed the inter-rater agreement, obtaining a Kendall Tau correlation coefficient of $0.34$. This translates into $69.8\%$ of all image pairs being concordant among annotators. Additionally, we performed statistical testing for the evaluations, and found that the voting results are statistically significant, at a p-value below 0.005.
13 Scalability
In the ablation study presented in Figure 3(d) of the main paper, we examine the visual appeal reward when we vary the numbers of generated images per prompt. Here, we provide a more detailed analysis of the extreme case based on 5 images per prompt, comparing Curriculum DPO with all the other fine-tuning strategies across all the three studied tasks. The results shown in Table 6 confirm that our method, Curriculum DPO, surpasses the competing methods even when the number of image samples per prompt is low. Therefore, we conclude that our training strategy does not require a high number of generated images per prompt to outperform DPO and DDPO.
14 Limitations
One limitation of our model is the introduction of additional hyperparameters, such as $B$ or $K$. These might require tuning in order to find the optimal values, which involves more computing power. However, in the ablation study from Section 4, we demonstrate that Curriculum DPO outperforms all baselines for multiple hyperparameter combinations. Therefore, suboptimal hyperparameter choices can still improve the generative models.
@@ 14 Limitations
15 Broader Impact
Generative models can be a valuable asset in many scenarios, ranging from boosting the productivity of creative tasks to being integrated in applications that are used on a daily basis (such as image restoration or super-resolution). Nevertheless, it might also represent a great source of fake data aiming for disinformation and impersonation, especially when the model is optimized to human preferences. In the recent years, an increase in deep fake materials flooded the Internet, with attackers aiming to spread false information or even steal sensitive information by posing as another entity or person.
While we strongly believe in the benefits of very capable generative models, we are aware of the potential risks. However, we can see that governments are working very closely with academia and industry on safely developing artificial intelligence, and thus observe and support the increasing focus on models that detect AI-generated content to mitigate the aforementioned risks. Notably, the ultimate goal of the project that funded our research is to develop robust deepfake detectors.